# Characteristics of drift effects in the quasi-geostrophic equation arising from nonlinear symmetry

### v1

@@ Abstract.
This paper compares two similar diffusion equations that appear in meteorology. One is the quasi-geostrophic equation, and the other is the convection-diffusion equation. Both are two-dimensional bilinear equations, and the order of differentiation is the same. Naturally, their scales also coincide. However, the direction in which the nonlinear effects act differs: one acts along the isothermal surface, while the other acts along the temperature gradient in a specified direction. The main assertion quantifies this difference through the large-time behavior of their solutions. In particular, the nonlinear distortions in the asymptotic profiles of both equations are compared. In this context, the spatial symmetry of the first approximation plays a crucial role, but the solutions require no symmetry.
@@ 1. Introduction
where $\partial_{t}=\partial/\partial t$ and $\mathcal{R}^{\bot}=(-\mathcal{R}^{2},\mathcal{R}^{1})$ for $\mathcal{R}^{j}\varphi=\partial_{j}(-\Delta)^{-1/2}\varphi=\mathcal{F}^{-1}[i\xi_{j}\mathcal{F}[\varphi]/|\xi|]$ is the orthogonal Riesz transform. The unknown function $\theta=\theta(t,x)\in\mathbb{R}$ describes potential temperature. This nonlinear term represents the effect of the Coriolis force $\mathcal{R}^{\bot}\theta$ acting along isothermal surfaces on the potential temperature. For this problem, global in time solvability in Lebesgue spaces is guaranteed by Nash theory. Moreover, the solution satisfies
@@
where $M_{0}=\int_{\mathbb{R}^{2}}\theta_{0}(x)dx=\int_{\mathbb{R}^{2}}\theta(t,x)dx$ is the conserved quantity and $G(t)=(4\pi t)^{-1}e^{-|x|^{2}/4t}$ is the fundamental solution (see[18]). Based on considerations of similar models, it is expected that this difference contains nonlinear components. We introduce the following convection-diffusion equation as a similar model:
@@
$\lambda^{2+1}(M_{1}\cdot\nabla G,J_{1})(\lambda^{2}t,\lambda x)=(M_{1}\cdot\nabla G,J_{1})(t,x)$
@@
for $t>0$. Both of $M_{1}\cdot\nabla G$ and $J_{1}$ are odd-type functions. However, any terms of $J_{1}$ exhibit higher symmetries than $M_{1}\cdot\nabla G$. Generally, in linear phenomena, based on the second law of thermodynamics, there is a clear correspondence between scale and symmetry. Nonlinear distortions are what break this correspondence. The convection diffusion is affected by the nonlinear distortion whenever $a\neq 0$. Another point to note is that the coefficients in the asymptotic expansion contain the ambiguous component $\int_{0}^{\infty}\int_{\mathbb{R}^{2}}((|\rho|\rho)(s,y)-|M_{0}|M_{0}G^{2}(1+s,y))dyds$. For the Burgers-type equation with the nonlinear term replaced by $a\cdot\nabla(\rho^{2})$ instead of $a\cdot\nabla(|\rho|\rho)$, we see that
@@
as $t\to+\infty$ for $1\leq q\leq\infty$ and $\gamma_{q}=1-\frac{1}{q}$, where $M_{0}=\int_{\mathbb{R}^{2}}\theta_{0}(x)dx\in\mathbb{R}$ and $M_{1}=-\int_{\mathbb{R}^{2}}x\theta_{0}(x)dx\in\mathbb{R}^{2}$. In addition, if $|x|^{2}\rho_{0}\in L^{1}(\mathbb{R}^{2})$, then the left-hand side is estimated by $O(t^{-\gamma_{q}-1}\log t)$ as $t\to\infty$.
That is, the second prediction is true and no effect of the nonlinear drift is appearing in the profile. Particularly, nonlinear distortions never be seen. Furthermore, neither logarithmic shift nor coefficient ambiguity is observed. The logarithmic growth in the sharp estimate is mitigated. This result is natural from a meteorological perspective. According to the linear diffution, the potential temperature asymptotically converges to the radially symmetric function as (1.4). Therefore, the Coriolis force acts most strongly on the radially symmetric component. This generates vortices due to geostrophic winds, but these vortices do not contribute to the averaging of potential temperature. Consequently, the nonlinear external forces do not affect the behavior of the solution.
The quasi-geostrophic equation was formulated as a dissipative model by Constantin, Majda and Tabak [4]. Subsequently, one with $(-\Delta)^{\alpha/2}$ instead of $\Delta$ was introduced. In particular, the case $\alpha=1$ is important as a critical phenomena in the abstract theory (cf.[2, 3, 5, 13, 15, 17]). However, in the context of asymptotic analysis, there is no critical case in $\alpha\leq 2$ since the scale critical is $\alpha=3/2$ (see[6, 7, 20]). As clearly expected, when $\alpha<2$, the nonlinear effects become even weaker than those in our problem.
@@
Notation. We often omit the spatial parameter from functions, for example, $u(t)=u(t,x)$. In particular, $G(t)*u_{0}=\int_{\mathbb{R}^{2}}G(t,x-y)u_{0}(y)dy$ and $\int_{0}^{t}g(t-s)*f(s)ds=\int_{0}^{t}\int_{\mathbb{R}^{2}}g(t-s,x-y)f(s,y)dyds$. We define the Fourier transform and its inverse by $\mathcal{F}[\varphi](\xi)=(2\pi)^{-1}$ $\int_{\mathbb{R}^{2}}e^{-ix\cdot\xi}\varphi(x)dx$ and $\mathcal{F}^{-1}[\varphi](x)=(2\pi)^{-1}\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}\varphi(\xi)d\xi$, where $i=\sqrt{-1}$. We denote the derivations by $\partial_{t}=\partial/\partial t,~\partial_{j}=\partial/\partial x_{j}~(j=1,2),~\nabla=(\partial_{1},\partial_{2}),~\nabla^{\bot}=(-\partial_{2},\partial_{1})$ and $\Delta=\partial_{1}^{2}+\partial_{2}^{2}$. The Riesz transform is defined by $R_{j}\varphi=\partial_{j}(-\Delta)^{-1/2}\varphi=\mathcal{F}^{-1}[i\xi_{j}|\xi|^{-1}\mathcal{F}[\varphi]]$ for $j=1,2$, and $\mathcal{R}^{\bot}=(-\mathcal{R}_{2},\mathcal{R}_{1})$. For $\beta=(\beta_{1},\beta_{2})\in\mathbb{Z}_{+}^{2}=(\mathbb{N}\cup\{0\})^{2},~|\beta|=\beta_{1}+\beta_{2}$. The Lebesgue space and its norm are denoted by $L^{q}(\mathbb{R}^{2})$ and $\|\cdot\|_{L^{q}(\mathbb{R}^{2})}$, that is, $\|f\|_{L^{q}(\mathbb{R}^{2})}=(\int_{\mathbb{R}^{2}}|f(x)|^{q}dx)^{1/q}$ for $1\leq q<\infty$ and $\|f\|_{L^{\infty}(\mathbb{R}^{2})}$ is the essential supremum. The heat kernel and its decay rate on $L^{q}(\mathbb{R}^{2})$ are symbolized by $G(t,x)=(4\pi t)^{-1}e^{-|x|^{2}/(4t)}$ and $\gamma_{q}=1-\frac{1}{q}$. We employ Landau symbol. Namely, $f(t)=o(t^{-\mu})$ and $g(t)=O(t^{-\mu})$ mean $t^{\mu}f(t)\to 0$ and $t^{\mu}g(t)\to c$ for some $c\in\mathbb{R}$ such as $t\to+\infty$ or $t\to+0$, respectively. Various nonnegative constants are denoted by $C$.
@@ 2.3. The case of quasi-geostrophic equation
$\begin{split}&e^{-r}=\frac{1}{\pi}\int_{\mathbb{R}}\frac{e^{-irx}}{1+x^{2}}dx=\frac{2\sqrt{\pi}}{\pi^{2}}\int_{\mathbb{R}}\int_{0}^{\infty}e^{-\lambda(1+x^{2})}d\lambda e^{-irx}dx\\ &=\frac{2\sqrt{2}}{\pi}\int_{0}^{\infty}e^{-\lambda}(4\pi\lambda)^{-1/2}e^{-r^{2}/(4\lambda)}d\lambda=\frac{2\sqrt{2}}{\pi^{2}}\int_{0}^{\infty}\lambda^{-3/2}e^{-4/\lambda}e^{-\lambda r^{2}}d\lambda.\end{split}$
@@
$\frac{i\xi^{\bot}}{|\xi|}=\frac{2\sqrt{2\pi}}{\pi^{3}}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-4/\lambda}i\xi^{\bot}e^{-\sigma^{2}\lambda|\xi|^{2}}d\lambda d\sigma.$
Consequently,
$\mathcal{R}^{\bot}G(t)=\frac{4}{\pi^{3}}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-4/\lambda}\nabla^{\bot}G(t+\sigma^{2}\lambda)d\lambda d\sigma.$
@@
since $\nabla^{\bot}\cdot\nabla G=0$. The ambiguous component $\int_{0}^{\infty}\int_{\mathbb{R}^{2}}((\theta\mathcal{R}^{\bot}\theta)(s,y)-M_{0}^{2}(G\mathcal{R}^{\bot}G)(1+s,y))dyds$ on $M_{1}$ is vanishing since $\mathcal{R}^{\bot}$ is skew-adjoint in $L^{2}(\mathbb{R}^{2})$. The logarithmic growth on the sharp estimate is mitigated. Indeed, since the logarithmic shift is omitted from (2.2), we see that $\|\theta(t)-M_{0}G(t)\|_{L^{q}(\mathbb{R}^{2})}=O(t^{-\gamma_{q}-\frac{1}{2}})$ as $t\to+\infty$. Employing this instead of (1.4) yields that

### v2

# Characteristics of drift effects arising from nonlinear symmetry of the quasi-geostrophic equation

@@ Abstract.
This paper compares two similar diffusion equations that appear in meteorology. One is the quasi-geostrophic equation, and the other is the convection-diffusion equation. Both are two-dimensional bilinear equations, and the order of differentiation is the same. Naturally, their scales also coincide. However, the direction in which the nonlinear effects act differs: one acts along the isothermal surface, while the other acts along the temperature gradient in a specified direction. The main assertion quantifies this difference through the large-time behavior of their solutions. In particular, the nonlinear distortions in the asymptotic profiles of both equations are compared. In this context, the spatial symmetry of the first approximation plays a crucial role, but the solutions require no symmetry. As an appendix, the mixed problem of those models are studied.
@@ 1. Introduction
where $\partial_{t}=\partial/\partial t$ and $\mathcal{R}^{\bot}=(-\mathcal{R}^{2},\mathcal{R}^{1})$ for $\mathcal{R}^{j}\varphi=\partial_{j}(-\Delta)^{-1/2}\varphi=\mathcal{F}^{-1}[i\xi_{j}\mathcal{F}[\varphi]/|\xi|]$ is the ‘orthogonal’ Riesz transform. The unknown function $\theta=\theta(t,x)\in\mathbb{R}$ describes potential temperature. This nonlinear term represents the effect of the Coriolis force $\mathcal{R}^{\bot}\theta$ acting along isothermal surfaces on the potential temperature. For this problem, global in time solvability in Lebesgue spaces is guaranteed by Moser–Nash theory. Moreover, the solution satisfies
@@
where $M_{0}=\int_{\mathbb{R}^{2}}\theta_{0}(x)dx=\int_{\mathbb{R}^{2}}\theta(t,x)dx$ is the conserved quantity and $G(t)=(4\pi t)^{-1}e^{-|x|^{2}/4t}$ is the fundamental solution (see[18]). Based on considerations of similar models, it is expected that the difference on (1.4) contains nonlinear components. We introduce the following convection-diffusion equation as a similar model:
@@
(1.7) $\lambda^{2+1}(M_{1}\cdot\nabla G,J_{1})(\lambda^{2}t,\lambda x)=(M_{1}\cdot\nabla G,J_{1})(t,x)$
@@
for $t>0$. Both of $M_{1}\cdot\nabla G$ and $J_{1}$ are odd-type functions. However, any terms of $J_{1}$ exhibit higher symmetries than $M_{1}\cdot\nabla G$. Generally, in linear phenomena, based on the second law of thermodynamics, there is a clear correspondence between scale and symmetry. Nonlinear distortions are what break this correspondence. The convection diffusion is affected by the nonlinear distortion whenever $a\neq 0$ since $\{a\cdot\nabla(-\Delta)^{k}G\}_{k\in\mathbb{N}}$ are independent. Another point to note is that the coefficients in the asymptotic expansion contain the ambiguous component $\int_{0}^{\infty}\int_{\mathbb{R}^{2}}((|\rho|\rho)(s,y)-|M_{0}|M_{0}G^{2}(1+s,y))dyds$. For the Burgers-type equation with the nonlinear term replaced by $a\cdot\nabla(\rho^{2})$ instead of $a\cdot\nabla(|\rho|\rho)$, we see that
@@
as $t\to+\infty$ for $1\leq q\leq\infty$ and $\gamma_{q}=1-\frac{1}{q}$, where $M_{0}=\int_{\mathbb{R}^{2}}\theta_{0}(x)dx\in\mathbb{R}$ and $M_{1}=-\int_{\mathbb{R}^{2}}x\theta_{0}(x)dx\in\mathbb{R}^{2}$. In addition, if $|x|^{2}\theta_{0}\in L^{1}(\mathbb{R}^{2})$, then the left-hand side is estimated by $O(t^{-\gamma_{q}-1}\log t)$ as $t\to\infty$.
That is, the second prediction is true and no effect of the nonlinear drift is appearing in the profile. Particularly, nonlinear distortions never be seen. Furthermore, neither logarithmic shift nor coefficient ambiguity is observed. The logarithmic growth in the sharp estimate is mitigated. This result is natural from a meteorological perspective. According to the linear diffution, the potential temperature asymptotically converges to the radially symmetric function as (1.4). We note that the initial conditions $x\theta_{0},~|x|^{2}\theta_{0}\in L^{1}(\mathbb{R}^{2})$ eliminate troughs and ridges. Therefore, the Coriolis force acts most strongly on the radially symmetric component. This generates smooth rotation due to geostrophic winds, but such a rotation contributes little to the averaging of potential temperature. Consequently, the nonlinear external forces do not affect the behavior of the solution. What is important in (1.1) is not that it contains the Riesz transform, but rather the direction in which it acts (see Appendix A).
The quasi-geostrophic equation was formulated as a dissipative model by Constantin, Majda and Tabak [4]. Subsequently, one with $(-\Delta)^{\alpha/2}$ instead of $\Delta$ was introduced. In particular, the case $\alpha=1$ is important as a critical phenomena in the abstract theory (cf.[2, 3, 5, 13, 15, 17]). However, in the context of asymptotic analysis, there is no critical case in $\alpha\leq 2$ since the scale critical is $\alpha=3$ (see[6, 7, 20]). As clearly expected, when $\alpha<2$, the nonlinear effects become even weaker than those in our problem.
@@
Notation. We often omit the spatial parameter from functions, for example, $u(t)=u(t,x)$. In particular, $G(t)*u_{0}=\int_{\mathbb{R}^{2}}G(t,x-y)u_{0}(y)dy$ and $\int_{0}^{t}g(t-s)*f(s)ds=\int_{0}^{t}\int_{\mathbb{R}^{2}}g(t-s,x-y)f(s,y)dyds$. We define the Fourier transform and its inverse by $\mathcal{F}[\varphi](\xi)=(2\pi)^{-1}$ $\int_{\mathbb{R}^{2}}e^{-ix\cdot\xi}\varphi(x)dx$ and $\mathcal{F}^{-1}[\varphi](x)=(2\pi)^{-1}\int_{\mathbb{R}^{2}}e^{ix\cdot\xi}\varphi(\xi)d\xi$, where $i=\sqrt{-1}$. We denote the derivations by $\partial_{t}=\partial/\partial t,~\partial_{j}=\partial/\partial x_{j}$ for $j=1,2,~\nabla=(\partial_{1},\partial_{2}),~\nabla^{\bot}=(-\partial_{2},\partial_{1})$ and $\Delta=\partial_{1}^{2}+\partial_{2}^{2}$. The Riesz transform is defined by $R_{j}\varphi=\partial_{j}(-\Delta)^{-1/2}\varphi=\mathcal{F}^{-1}[i\xi_{j}|\xi|^{-1}\mathcal{F}[\varphi]]$ for $j=1,2$, and $\mathcal{R}^{\bot}=(-\mathcal{R}_{2},\mathcal{R}_{1})$ and $\mathcal{R}=(\mathcal{R}_{1},\mathcal{R}_{2})$. For $\beta=(\beta_{1},\beta_{2})\in\mathbb{Z}_{+}^{2}=(\mathbb{N}\cup\{0\})^{2},~|\beta|=\beta_{1}+\beta_{2}$. The Lebesgue space and its norm are denoted by $L^{q}(\mathbb{R}^{2})$ and $\|\cdot\|_{L^{q}(\mathbb{R}^{2})}$, that is, $\|f\|_{L^{q}(\mathbb{R}^{2})}=(\int_{\mathbb{R}^{2}}|f(x)|^{q}dx)^{1/q}$ for $1\leq q<\infty$ and $\|f\|_{L^{\infty}(\mathbb{R}^{2})}$ is the essential supremum. The heat kernel and its decay rate on $L^{q}(\mathbb{R}^{2})$ are symbolized by $G(t,x)=(4\pi t)^{-1}e^{-|x|^{2}/(4t)}$ and $\gamma_{q}=1-\frac{1}{q}$. We employ Landau symbol. Namely, $f(t)=o(t^{-\mu})$ and $g(t)=O(t^{-\mu})$ mean $t^{\mu}f(t)\to 0$ and $t^{\mu}g(t)\to c$ for some $c\in\mathbb{R}$ such as $t\to+\infty$ or $t\to+0$, respectively. Various nonnegative constants are denoted by $C$.
@@ 2.3. The case of quasi-geostrophic equation
$\begin{split}&e^{-r}=\frac{1}{\pi}\int_{\mathbb{R}}\frac{e^{-irx}}{1+x^{2}}dx=\frac{2\sqrt{\pi}}{\pi^{2}}\int_{\mathbb{R}}\int_{0}^{\infty}e^{-\lambda(1+x^{2})}d\lambda e^{-irx}dx\\ &=\frac{4\sqrt{\pi}}{\pi}\int_{0}^{\infty}e^{-\lambda}(4\pi\lambda)^{-1/2}e^{-r^{2}/(4\lambda)}d\lambda=\frac{1}{\pi}\int_{0}^{\infty}\lambda^{-3/2}e^{-\frac{1}{4\lambda}}e^{-\lambda r^{2}}d\lambda.\end{split}$
@@
$\mathcal{R}^{\bot}G(t)=\frac{4\sqrt{\pi}}{\pi}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-\frac{1}{4\lambda}}\nabla^{\bot}G(t+\sigma^{2}\lambda)d\lambda d\sigma.$
@@
since $\nabla^{\bot}\cdot\nabla G=0$. The ambiguous component $\int_{0}^{\infty}\int_{\mathbb{R}^{2}}((\theta\mathcal{R}^{\bot}\theta)(s,y)-M_{0}^{2}(G\mathcal{R}^{\bot}G)(1+s,y))dyds$ on $M_{1}$ is vanishing since $\mathcal{R}^{\bot}$ is skew-adjoint in $L^{2}(\mathbb{R}^{2})$. The logarithmic growth on the sharp estimate is mitigated. Indeed, since the logarithmic shift is omitted from (2.3), we see that
(2.5) $\|\theta(t)-M_{0}G(t)\|_{L^{q}(\mathbb{R}^{2})}=O(t^{-\gamma_{q}-\frac{1}{2}})$
as $t\to+\infty$. Employing this instead of (1.4) yields that
@@
Appendix A A force acting perpendicular to the isothermal surface and equal in effect to the Coriolis force
Here, to see how uniquely the Coriolis force behaves in (1.1), we introduce the following artificial problem:
(A.1) $\left\{\begin{array}[]{lr}\partial_{t}\vartheta-\nabla\cdot(\vartheta\mathcal{R}\vartheta)=\Delta\vartheta,&t>0,~x\in\mathbb{R}^{2},\\ \vartheta(0,x)=\vartheta_{0}(x),&x\in\mathbb{R}^{2},\end{array}\right$.
where $\mathcal{R}=(\mathcal{R}^{1},\mathcal{R}^{2})$ is the ‘forward-directed’ Riesz transform. This equation exhibits a nonlinear external force equivalent to the Coriolis force. However, unlike the quasi-geostrophic equation, this force acts perpendicular to the isothermal surface. In this sense, (A.1) possesses a symmetry distinct also from that of the convection-diffusion equation. Under the $L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$-framework, Moser–Nash theory guarantees global in time existence of solutions and (1.2) for (A.1). Indeed,
$\int_{\mathbb{R}^{2}}|\vartheta|^{q-2}\vartheta\nabla\cdot(\vartheta\mathcal{R}\vartheta)dx=\frac{q-1}{q}\int_{\mathbb{R}^{2}}\mathcal{R}\vartheta\cdot\nabla(|\vartheta|^{q})dx=\frac{q-1}{q}\int_{\mathbb{R}^{2}}|\vartheta|^{q}\Lambda\vartheta dx$
for $\Lambda\vartheta=\mathcal{F}^{-1}[|\xi|\hat{\vartheta}$] and, hence, Stroock–Varopoulos estimate treats this term (cf.[7]). Now we recall Section 2.1. The solution $\vartheta$ satisfies (2.1) for $(u_{0},\boldsymbol{f})=(\vartheta_{0},\vartheta\mathcal{R}\vartheta)$. Then, the weighted estimate (1.3) is derived from the $L^{p}$-$L^{q}$ estimate together with (1.2). The logarithmic shift
$\int_{0}^{t}\int_{\mathbb{R}^{2}}\boldsymbol{f}[M_{0}G](1+s,y)dyds=M_{0}^{2}\int_{0}^{t}(1+s)^{-1}ds\int_{\mathbb{R}^{2}}(G\mathcal{R}G)(1,y)dy$
is vanishing since $G\mathcal{R}G$ is also odd-type. Similarly, the unclear components within $M_{1}$ disappear. On the other hand, the nonlinear distortion appears. Indeed, on the same way as in Section 2.3, we see that
$\begin{split}&J_{1}(t)=M_{0}^{2}\int_{0}^{t}\nabla G(t-s)*(G\mathcal{R}G)(s)ds\\ &=\frac{4\sqrt{\pi}M_{0}^{2}}{\pi}\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-\frac{1}{4\lambda}}\nabla G(t-s)*(G(s)\nabla G(s+\sigma^{2}\lambda))d\lambda d\sigma ds\\ &=\frac{\sqrt{\pi}M_{0}^{2}}{\pi^{2}}\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-\frac{1}{4\lambda}}s(2s+\sigma^{2}\lambda)^{-2}\Delta G(t-\tfrac{s^{2}}{2s+\sigma^{2}\lambda})d\lambda d\sigma ds.\end{split}$
Here, $t-\frac{s^{2}}{2s+\sigma^{2}\lambda}\geq\frac{t}{2}$ for $(\lambda,\sigma,s)\in(0,\infty)^{2}\times(0,t)$, then $\Delta G(t-\tfrac{s^{2}}{2s+\sigma^{2}\lambda})$ is analytic around $\tfrac{s^{2}}{2s+\sigma^{2}\lambda}=0$ and we see from the elemental calculus that
$\begin{split}&J_{1}(t)=\frac{\sqrt{\pi}M_{0}^{2}}{\pi^{2}}\sum_{k=0}^{\infty}\frac{(-\Delta)^{k+1}G(t)}{k!}\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\infty}\lambda^{-3/2}e^{-\frac{1}{4\lambda}}s^{-3}\left(\frac{s^{2}}{2s+\sigma^{2}\lambda}\right)^{k+2}d\lambda d\sigma ds\\ &=\frac{\sqrt{2\pi}M_{0}^{2}}{2\pi}\sum_{k=0}^{\infty}\left(\frac{t}{2}\right)^{k+\frac{1}{2}}\frac{(2k-1)!!(-\Delta)^{k+1}G(t)}{k!(2k+2)!!}.\end{split}$
Remember that any terms in this function have the same scale, that is, (1.7) holds also for this $J_{1}$. From the same reason as in Section 2.3, we see that $\|r_{2}(t)\|_{L^{q}(\mathbb{R}^{2})}=O(t^{-\gamma_{q}}\log t)$ as $t\to+\infty$. Summarizing the above, we see the following result.
Proposition A.1.
Let $\vartheta_{0}\in L^{1}(\mathbb{R}^{2})\cap L^{\infty}(\mathbb{R}^{2})$ and $x\vartheta_{0}\in L^{1}(\mathbb{R}^{2})$. Then the solution $\vartheta$ of (A.1) fulfills that
$\|\vartheta(t)-M_{0}G(t)-M_{1}\cdot\nabla G(t)-J_{1}(t)\|_{L^{q}(\mathbb{R}^{2})}=o(t^{-\gamma_{q}-1/2})$
as $t\to+\infty$ for the above $J_{1}$, and $1\leq q\leq\infty$ and $\gamma_{q}=1-\frac{1}{q}$, where $M_{0}=\int_{\mathbb{R}^{2}}\vartheta_{0}(x)dx\in\mathbb{R}$ and $M_{1}=-\int_{\mathbb{R}^{2}}x\vartheta_{0}(x)dx\in\mathbb{R}^{2}$. In addition, if $|x|^{2}\vartheta_{0}\in L^{1}(\mathbb{R}^{2})$, then the left-hand side is estimated by $O(t^{-\gamma_{q}-1}\log t)$ as $t\to\infty$.
Compared to the previous results, unlike the convection-diffusion equation, no logarithmic shift nor ambiguous coefficients appear in the solution behavior. From this perspective, the nonlinearity of (A.1) resembles that of the quasi-geostrophic equation. On the other hand, unlike the quasi-geostrophic equation, the nonlinear distortion $J_{1}$ appears. However, this distortion possesses a symmetry distinct from that of the convection-diffusion equation. Precisely, $J_{1}$ for (A.1) is radially symmetric and the derivative order is shifted. One final comment: like the convection-diffusion equation, this distortion reflects the nonlinear structure of (A.1), i.e., $(-\Delta)^{k+1}G=-(-\Delta)^{k+1/2}\nabla\cdot\mathcal{R}G$.